OpenStudy (anonymous):

Find the solution of log x + log (x+1) = log 12.

OpenStudy (anonymous):

Hint: \(\log(mn) = \log(m)+\log(n)\)

OpenStudy (xapproachesinfinity):

use the product rule for logs \[\log a+\log b= \log\ ab\]

OpenStudy (anonymous):

Yeah, I already got log(x^2+x) = log 12, but what would I do after that?

OpenStudy (anonymous):

well, you basically can cancel log out. \(x^2+x\) must be equal to 12. Make sense?

OpenStudy (anonymous):

So I can "legally" cancel out logs with the same bases all of the time?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Great, thanks!